find 2 consecutive integers whos sum is 279: find 2 consecutive integers whos sum is 279
Answer by josmiceli(2035)

The consecutive integers can be called n and n+1
n + (n + 1) = 279
2n + 1 = 279
2n = 278
n = 139
n + 1 = 140
The consecutive integers are 139 and 140